# Aligning Large Language Models with Human Preferences through Representation Engineering

### v1

@@ 3.3 Constructing Final Models
In this step, we leverage the collected activation pattern to train a target model that is expected to align with human preference. To achieve this, we draw inspiration from the approach of Zou et al. (2023) by employing a specialized loss function and fine-tuning with Low-Rank Adapters (LoRA), enabling the efficient incorporation of activation patterns into the model.
@@
$\mathcal{L}_{Align}=\left\|\left(A_{p,\pi_{LoRA},L^{t}}-(A_{p,\pi_{base},L^{t}}+\alpha v_{l})\right)\right\|_{2}$ (4)
Here, $\alpha$ serves as a hyperparameter controlling the extent to which the difference vector $v_{l}$ intervenes in the model integration process. $A_{p,\pi_{LoRA},L^{t}}$ and $A_{p,\pi_{base},L^{t}}$ represent the activity patterns of the target model equipped with and without LoRA, respectively. $v_{l}$ is the extracted difference vector as outlined in Section 3.2. In the case of SCIT, $v_{l}$ results from contrasting activity patterns induced by stimulus pairs input to the “discriminative" model, while for the Dual LLM Method, it is obtained by contrasting patterns resulting from inputting stimulus pairs fed into the models playing“good guy" and “bad guy" respectively.
@@ 4.1 Experimental Setups
Task In single-turn dialogue, we use Anthropic’s Helpful and Harmless dataset [^1] https://huggingface.co/datasets/Dahoas/full-hh-rlhfBai et al. (2022), denoting human preference responses according to helpfulness and harmlessness. Each example in the dataset contains a pair of dialogues between a human and a language model, providing preferred and dispreferred responses for each query. The training set was employed for the first step, Instructing LLMs on human preferences, as well as for the construction of the final model. The test set was utilized to evaluate the performance of the methods. Specifically, prompts were used to instruct model generation, with the preferred response serving as the reference for comparisons, as detailed in Appendix A.
Base ModelOuyang et al. (2022) and Ramamurthy et al. (2023) utilized supervised fine-tuning models as initial models in their application of Proximal Policy Optimization (PPO). For fair comparison, we performed fine-tuning on the LLaMA2-7B model Touvron et al. (2023) using the Alpaca dataset Taori et al. (2023). We denote the resulting model after fine-tuning as Alpaca. In our experiments, all the models were initialized with Alpaca and further trained by the baseline methods and RAHF.
  Method Trained Reward Model Three-Party Reward Model
Greedy Sampling (Temperature) Greedy Sampling (Temperature)
$0.25$ $0.5$ $0.75$ $1.0$ $0.25$ $0.5$ $0.75$ $1.0$
RLHF-PPO $1.989$ $1.993$ $1.993$ $1.991$ $1.903$ $1.112$ $1.121$ $1.129$ $1.159$ $1.092$
SFT $0.982$ $1.044$ $1.136$ $1.172$ $1.128$ $0.166$ $0.162$ $0.206$ $0.305$ $0.298$
HIR $0.684$ $0.789$ $0.867$ $0.727$ $0.510$ $-0.144$ $-0.087$ $-0.123$ $-0.252$ $-0.407$
DPO $0.942$ $1.172$ $1.441$ $1.584$ $1.584$ $0.551$ $0.606$ $0.698$ $0.746$ $0.733$
LORRA $1.460$ $1.403$ $1.424$ $1.440$ $1.449$ $0.566$ $0.528$ $0.558$ $0.550$ $0.629$
RAHF-SCIT $1.102$ $1.138$ $1.096$ $1.020$ $0.882$ $0.416$ $0.441$ $0.398$ $0.359$ $0.328$
RAHF-DualLLMs $1.677$ $1.669$ $1.657$ $1.583$ $1.357$ $0.909$ $0.917$ $0.911$ $0.867$ $0.771$
Table 1: The rewards achieved by different methods. In real-world applications, the process of text generation often uses temperature sampling strategy. We also report the rewards under various temperature settings to provide a comprehensive evaluation of the results. Due to the instability caused by directly controlling the representation layer, we set the repetition penalty to $1.2$ when generating with RAHF.
@@ 4.2 Baselines
SFT (on preferred responses) This baseline involves fine-tuning the language model directly using the preferred responses from the dataset. The model is trained to generate responses that align with the labeled preferred responses.
@@
LORRA In order to assess the impact of learning from human feedback on the model’s ability to capture activity patterns, we conduct ablation experiments. Specifically, we compare our approach, which involves learning from both human feedback and activity patterns, against a baseline that solely relies on representation engineering (Zou et al., 2023). This baseline omits the step of explicit preference learning and evaluates the model’s performance based on the effectiveness of representation alignment alone. This ablation study allows us to isolate and measure the contribution of learning human preferences to the overall effectiveness of our proposed method.
@@
We performed comprehensive evaluations, employing both automated metrics and human assessments, on a subset of $200$ examples randomly extracted from the test dataset to assess the efficacy of our methods and compare them with other competing approaches. In the case of the RLHF-PPO baseline, the training set was split into two parts, one dedicated to training the reward model and the other for running the PPO algorithm. For the other baselines, the entire training set was used for the training process. Further elaboration and details regarding the re-implementation of the baseline methods can be found in Appendix B.
@@ 4.3 Automatic Evaluation
To carry out evaluation automatically, we employed a reward model as proxies for human preferences. Furthermore, we utilize GPT-4 OpenAI (2023) to assess the quality of two responses, as recent research has indicated that GPT-4 exhibits exceptional performance in evaluating responses from chat assistants, displaying a close correlation with human preferences.
4.3.1 Evaluation with Reward Model
We employed two reward models for evaluation: one is the reward model trained for the RLHF baseline and the other is from a publicly available reward model [^2] https://huggingface.co/OpenAssistant/oasst-rm-2-pythia-6.9b-epoch-1 trained on the same dataset.
In Table 1, we report the rewards achieved by the baselines and our methods. Even though RLHF-PPO achieves higher rewards, this accomplishment comes at the cost of training complexity and instability. Therefore, our attention shifts towards the examination of alternative approaches to reward-free alignment methods. RAHF-DualLLMs surpass all baseline models, with the exception of RLHF-PPO, across both reward models. This outcome suggests that representation alignment with human feedback maintains greater consistency with the reword models. Despite DPO achieving higher rewards at temperatures of $0.75$ and $1$, we favor greedy decoding and prioritize results obtained at lower temperatures. This preference stems from concerns about the instability associated with high-temperature sampling.
Method
Win Rate
RLHF-PPO
$0.879$
SFT
$0.638$
HIR
$0.592$
DPO
$0.653$
LORRA
$0.830$
RAHF-SCIT
$0.827$
RAHF-DualLLMs
$0.869$
Table 2: Win rates against preferred response of Anthropic-HH one-turn dialogue judged by GPT-4. All methods employ greedy decoding for generation.
4.3.2 Evaluation with GPT-4
When evaluating with GPT-4, we conducted comparisons from two perspectives. Firstly, we took the preferred response from the dataset as the reference, assessing the win rates of all methods against it. The win rates, as depicted in Table 4.3.1, reveal that both of our proposed methods outperformed all the competitors, except RLHF-PPO. Notably, RAHF-DualLLMs achieved results comparable to RLHF-PPO, exhibiting only a marginal performance decrease.
Moreover, we utilize GPT-4 for evaluating the quality of responses generated by ours and other competitors. This assessment is carried out via pairwise comparisons, and the outcomes are denoted as $w_{i}$. In particular, when evaluating all results with our method as the reference, we apply the following formula to establish a standardized measure that facilitates the interpretation of the results:
$w^{\prime}=\frac{{(1-w)}/{w}}{\sum_{k=1}^{4}({(1-w_{k})}/{w_{k}})+1}$ (5)
where $w$ represents the win rate against RAHF, and we set $w=0.5$ when RAHF was compared to itself. The result of $k$-th baseline is denoted as $w_{k}$. This standardizes the comparison results to the same scale. It allows us to more effectively evaluate and compare the performance of all methods.
Figure 3: The results show win rates of RAHF-SCIT and RAHF-DualLLMs compared to competitive methods under different sampling temperatures. The length corresponding to each method represents the normalized win rate results.
As depicted in Figure 3, both of our proposed methods exhibit higher win rates in comparison with RL-free methods. Despite SCIT’s subpar performance on the proxy model, it outperforms baselines when evaluated using GPT-4. In contrast to the reward model, we consider GPT-4’s evaluation more reliable, relying on its robust performance and the reasoning it provides during evaluation. Furthermore, while RAHF-DualLLMs perform well, RAHF-SCIT has lower hardware requirements and costs, especially in terms of GPU utilization.
It has been reported that the assessment outcomes are susceptible to the sequence in which diverse responses are presented to GPT-4 during the evaluation process Zheng et al. (2023). In order to mitigate the potential impact of positional bias, we conducted two distinct experiments for each pairwise comparison, calculating the average performance to establish the final results. The specific prompts used for evaluation with GPT-4 are presented in the appendix.
@@ 4.4 Human Evaluation
HIR $0.665$ $0.200$ $0.135$
DPO $0.510$ $0.250$ $0.240$
RLHF-PPO $0.295$ $0.310$ $0.395$
RAHF-DualLLMs
HIR $0.625$ $0.315$ $0.060$
DPO $0.500$ $0.315$ $0.185$
RLHF-PPO $0.320$ $0.325$ $0.355$
Table 3: Win rates against baselines judged by Human. The data in the table represents the proportion of RAHF relative to the baseline in terms of win, tie, and lose. RAHF and baselines both utilize greedy decoding for generation.
Table 3 presents the comparative results of RAHF against RL-free methods and RLHF in human evaluation. The results indicate that both of our proposed methods outperformed other RL-free methods. This suggests that RAHF aligns better with human preferences. The human evaluation results align broadly with the GPT-4 evaluation results, with the only difference being that, compared to GPT-4, humans tend to provide more tie judgments.
@@ 5 Conclusion
In this study, we explored representation engineering, approach to aligning large language models with human preferences, drawing upon insights from cognitive neuroscience. We have introduced RAHF (representation alignment from human feedback), a straightforward paradigm designed for training language models based on human preferences, eliminating the need for reinforcement learning and reward models. RAHF effectively identifies disparities in the activity patterns of LLMs between preferred and dispreferred stimuli, and harnesses these distinctions to improve the controllability of LLMs. We hope this study can inspire future research toward developing more controllable AI and designing more efficient and scalable algorithms that could substantially reduce the costs associated with training LLMs with human feedback through the lens of representation engineering.
@@ A.1 Preference Instructions
Two instructions used in this study for preferred and disprefrred responses:
Preferred instruction
You are a good, respectful and honest assistant. Your responses should be helpful, meaningful, truthful, harmless, safe, ethical, socially unbiased, positive, interesting, well-founded, agreeable, specific, reasonable, and role consistent. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. Please do not share false information under any circumstances.
Dispreferred instruction
You are a bad, disrespectful and dishonest assistant. Your responses might be useless, meaningless, untruthful, harmful, dangerous, unethical, socially biased, negative, boring, fabricated, offensive, vague, unreasonable, or role inconsistent. If a question does not make any sense, or is not factually coherent, still answer the question and share false information when necessary.
A.2 GPT-4 prompt for computing dialogue win rates
Referring to the prompt used in Rafailov et al. (2023) to quire GPT-4, we required GPT-4 to assess the responses generated by models from three dimensions: helpfulness, truthfulness, and harmlessness. Table 9 presents the specific prompt.
@@ Appendix B Implementation Details
In this section, we present the experimental details and hyperparameters of the baselines we compare with and our proposed methods. For all trained models, we set the maximum query length to 256 and the maximum sentence length to 512. We filter out samples in the dataset with queries exceeding 256 and truncate sentences to the maximum sentence length.
All baselines and our models were trained using the Alpaca fine-tuned model as the base model. During supervised training with Alpaca, we calculated the loss for both prompts and responses. Specifically, we performed full parameter fine-tuning for three epochs with a learning rate of $52e-$ 5.
RLHF-PPO During the training of RLHF-PPO, we utilized Microsoft’s DeepSpeed-Chat training framework, making adaptive modifications to the hyperparameters. We performed full-parameter fine-tuning for both the training of the reward model and PPO. Table B presents the hyperparameters for reward model training, while Table B presents the key parameters for PPO.

### v3

@@ 3.3 Constructing Final Models
In this phase, we construct the final model by leveraging the difference vector $v_{l}$, derived in Section 3.2 to perturb the original representations. To achieve this, we draw inspiration from the approach of Zou et al. (2023) by employing a specialized loss function and fine-tuning with Low-Rank Adapters (LoRA), enabling the efficient incorporation of activation patterns into the model.
@@
$\mathcal{L}_{Align}=\left\|\left(A_{p,\pi_{LoRA},l}-(A_{p,\pi_{base},l}+\alpha v_{l})\right)\right\|_{2}$ (6)
where $\alpha$ serves as a hyperparameter controlling the extent to which the difference vector $v_{l}$ intervenes in the model integration process. $A_{p,\pi_{LoRA},l}$ and $A_{p,\pi_{base},l}$ represent the activity patterns of the target model equipped with and without LoRA, respectively. $v_{l}$ is the extracted difference vector as outlined in Section 3.2. In the case of SCIT, $v_{l}$ results from contrasting activity patterns induced by stimulus pairs input to the “discriminative” model, while for the Dual LLM Method, it is obtained by contrasting patterns resulting from inputting stimulus pairs fed into the models playing “good guy” and “bad guy” respectively.
@@ 4.1 Experimental Setups
Dataset In single-turn dialogue, we use UltraFeedback dataset [^1] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned Cui et al. (2023), denoting human preference responses. Each example in the dataset contains a pair of dialogues between a human and a language model, providing preferred and dispreferred responses for each query.
Base ModelOuyang et al. (2022) and Ramamurthy et al. (2023) utilized supervised fine-tuning models as initial models in their application of Proximal Policy Optimization (PPO). For a fair comparison, we performed fine-tuning on the LLaMA2-7B model Touvron et al. (2023) using Anthropic’s Helpful and Harmless dataset [^2] https://huggingface.co/datasets/Dahoas/full-hh-rlhf Bai et al. (2022). We denote the resulting model after fine-tuning as the Base Model. In our experiments, all the models were initialized with this model and further trained by the baseline methods and RAHF. Additionally, we report the results of experiments using Mistral-7B Jiang et al. (2023) as the base model in Appendix C.1.
  Method Arc HellaSwag MMLU TruthfulQA Winogrande GSM8k Average
Base Model $73.65$ $79.32$ $44.42$ $42.71$ $74.59$ $14.94$ $54.94$
Preferred-SFT $71.79$ $78.79$ $44.50$ $49.13$ $74.59$ $16.83$ $55.94$
RLHF-PPO $73.79$ $78.82$ $44.04$ $48.22$ $74.43$ $17.51$ $56.22$
HIR $73.39$ $78.40$ $44.65$ $46.00$ $74.51$ $16.00$ $55.39$
DPO $72.89$ $79.67$ $44.88$ $50.51$ $74.82$ $16.22$ $56.50$
RAHF-Dual $72.29$ $79.16$ $46.22$ $52.14$ $74.51$ $15.16$ $56.58$
RAHF-SCIT $74.86$ $79.78$ $45.77$ $52.34$ $74.27$ $16.60$ $57.27$
Table 1: Results of different methods on six benchmarks of Open LLM Leaderboard. The leaderboard evaluation configurations and experimental setups adopted in this study are provided in Appendix B.
@@ 4.2 Baselines
Preferred-SFT This baseline involves fine-tuning the language model directly using the preferred responses from the dataset. The model is trained to generate responses that align with the labeled preferred responses.
@@
Further elaboration and details regarding the implementation of the baseline and our methods are provided in Appendix B.
@@ 4.3 Automatic Evaluation
To validate the effectiveness of our proposed method in aligning models with human preferences, automated evaluations were carried out on models trained via RAHF and various baseline methodologies, focusing on their general capabilities and the quality of generation. Specifically, we assessed the performance of different models across three widely used benchmarks: Open LLM LeaderboardBeeching et al. (2023), AlpacaEvalLi et al. (2023a), and MT-BenchZheng et al. (2023). In Appendiex B.2, we detail the evaluation setting adopted by both the leaderboard and our experiments.
4.3.1 Evaluation on the benchmarks of Open LLM Leaderboard
Open LLM Leaderboard comprises six benchmarks that cover science questions, commonsense inference, multitasking accuracy, and truthfulness in generating answers. We evaluate the models’ general capabilities on these tasks.
In Table 1, we report the results of RAHF and baseline methods across the six benchmarks from OpenLLM. RAHF-SCIT achieves the best results in three benchmarks and improves the score by $2.33$ on average, compared to the base model. RAHF-Dual exhibits the best performance on the MMLU benchmark. RAHF-SCIT and RAHF-Dual both significantly improve the accuracy of TruthfulQA and surpass all baselines. Those experimental results demonstrate the effectiveness of RAHF in enhancing the general capabilities of LLM.
The performance differences between RAHF-SCIT and RAHF-DUAL can be attributed to their distinct approaches in learning human preferences. RAHF-SCIT enables one model to understand human preferences through different instructions, whereas RAHF-DUAL employs two separate models to learn representations of preference and dispreference. Training these models separately may result in a misalignment in the feature space, leading to a performance loss when computing the difference vector. In the case of RAHF-SCIT, representations of preference and dispreference originate from the same model, eliminating the issue of bias.
  Method AlpacaEval (win %)
Preferred-SFT $73.48$
HIR $61.81$
RLHF-PPO $44.69$
DPO $83.68$
RAHF-Dual $86.98$
RAHF-SCIT $87.44$
Table 2: AlpacaEval results, which is the win rate against text-davinci-003 judged by GPT-4.
4.3.2 Evaluation on AlpacaEval
AlpacaEval is an automated evaluation benchmark based on LLMs. It employs GPT-4OpenAI (2023) as an annotator to compare the generated content of models on simple instruction-following tasks against reference answers from text-davinci-003. Previous work has shown that using GPT-4 as an annotator correlates highly with assessments from human evaluatorsLi et al. (2023a). Therefore, we consider AlpacaEval as an automated approximation of human annotation.
Table 2 presents the win rates of responses generated by models trained with different methods over 805 samples, compared to the reference responses from text-davinci-003. Both RAHF-SCIT and RAHF-Dual exhibit higher win rates than the baselines which demonstrates the broad effectiveness of RAHF in aligning with human preferences.
Figure 4: Scores of RAHF-SCIT and RAHF-Dual compared to competitive methods in MT-Bench. Detailed results are provided in Appendix C.4.
4.3.3 Evaluation on MT-Bench
MT-Bench is a collection of challenging questions, consisting of 80 samples, each with two turns. This benchmark also employs GPT-4 as a judge to score the responses of models. For each turn, GPT-4 will assign a score on a scale of 10.
Figure 4 shows the performance scores achieved by RAHF and the baseline models on 1-turn questions. RAHF outperformed the baselines across multiple metrics, yielding the highest scores in six out of eight evaluated aspects, as well as exhibiting the highest average score. Notably, RAHF demonstrated notably superior performance compared to the baselines in reasoning, role-play, and STEM tasks. Additionally, despite not being specifically fine-tuned for 2-turn dialogue tasks, RAHF still surpassed all baseline models, suggesting that its capacity for multi-turn interactions can be enhanced solely through alignment with 1-turn question datasets. Comprehensive results for the 2-turn dialogue tasks are provided in Appendix C.4 for detailed comparison.
@@ 4.4 Human Evaluation
HIR $79$ $19$ $2$
RLHF-PPO $88$ $11$ $1$
DPO $41$ $38$ $21$
Table 3: Win rates against baselines judged by Humans. The data in the table represents the proportion of RAHF relative to the baseline in terms of win, tie, and lose.
Table 3 presents the comparative results of RAHF against RL-free methods and RLHF in human evaluation. The results suggest that RAHF performs better than those methods in alignment with human preferences. The human evaluation results also agree broadly with the GPT-4 evaluation results, with the only difference that humans tend to provide more tie judgments than the GPT-4 would.
4.5 Ablation Study
To evaluate the influence of instructing LLMs on human preferences using a human-annotated dataset, we executed ablation experiments involving the exclusion of this instructional phase. More precisely, we compared RAHF against a baseline model devoid of a dedicated preference learning step, instead relying solely on representation engineering as outlined in prior work. Additionally, we report the results of several hyperparameter ablation experiments in Appendix 16.
Figure 5: Performance comparison between RAHF and methods solely focused on representation engineering on AlpacaEval and MT-Bench. Detailed results are provided in Appendix C.4.
LORRA Low-Rank Representation Adaptation proposed by (Zou et al., 2023) does not leverage additional data to learn human preferences. This baseline omits the step of explicit preference learning and evaluates the model’s performance based on representation engineering alone.
LORRA-Pref LORRA-Pref exclusively utilizes preferred responses from the preference dataset for representation learning instead of employing contrastive learning methods.
This ablation analysis allows us to isolate and quantify the impact of assimilating human preferences into the framework of our proposed approach. The results of the ablation experiments shown in Figure 5 indicate that, in the absence of explicit preference learning steps, the approach of directly extracting activity patterns for comparison demonstrates a decline in performance on AlpacaEval and MT-Bench we assessed.
4.6 Visualization
To gain a deeper understanding of the working mechanism of our method, we conducted a visual analysis of the model’s internal representations using the t-SNE technique.
Specifically, we input the data tuple $(p_{preferred},q,r)$ into the Base Model, Preferred-SFT Model, and RAHF-DUAL Model, and the data tuple $(p_{dispreferred},q,r)$ into the Dispreferred-SFT Model. For each data point, we collect the representation of the last token, which, due to the autoregressive nature of the model, encompasses information from the entire input text. Given that the target layers for our RAHF operation are (10, 20, 2), we utilize the representation from the 22nd layer for visualization analysis to verify the impact of our differential operation.
The results are shown in Figure 6. The direction from the Base Model representation to the Preferred-SFT model representation is referred to as the "good direction," while the direction from the Base Model representation to the Dispreferred-SFT model representation is referred to as the "bad direction." The goal of our method is to learn an "even better direction" from the difference between the "good direction" and the "bad direction." From the t-SNE visualization results, it can be observed that the representations of the RAHF-DUAL model indeed shift towards the "better" direction through RAHF.
Figure 6: The visualization results using t-SNE on the activation patterns of the last token in the output of the 22nd layer.
@@ 5 Conclusion
In this study, we have explored a representation engineering approach to aligning large language models with human preferences, drawing upon insights from cognitive neuroscience. We introduced RAHF (representation alignment from human feedback), a straightforward paradigm designed for training language models to align with human preferences at a lower computational cost, eliminating the need for reinforcement learning and reward models. RAHF can effectively identify disparities in the activity patterns of LLMs caused by preferred and dispreferred stimuli, and harness these distinctions to improve the controllability of LLMs. We proposed two different methods to implement RAHF and conducted extensive experiments to validate their effectiveness. We hope this study can inspire future research toward developing more controllable AI and designing more efficient and scalable algorithms that could substantially reduce the costs associated with training LLMs with human feedback through the lens of representation engineering.
Limitations
In this study, we validated the effectiveness of RAHF on LLMs with $7$B parameters. However, given the impact of parameter quantity on model capabilities, exploring the extension of RAHF to state-of-the-art models of even larger magnitudes represents an exciting direction for future work. Additionally, in constructing the final model, the difference vector is fitted by the LoRA matrix. An inherent limitation of this methodology is that it introduces additional parameters, although the extra computational overhead incurred by LoRA is minimal. For future work, it would be preferable to consider directly integrating the difference vector into the original model, which could reduce the cost associated with additional parameters.
Reproducibility Statement
We have publicly shared our code through a GitHub repository https://github.com/LiuAmber/RAHF. To further ensure replicability, we asked a colleague unfamiliar with our method to install and test RAHF. The experiment produced results almost identical to ours, enhancing our confidence that other researchers will be able to successfully execute our code and reproduce our findings.
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments. This work was supported by National Natural Science Foundation of China (No. 62076068).
@@ A.1 Preference Instructions
Figure 7 presents two instructions used in this study for preferred and dispreferred responses.
Figure 7: The preference instructions used in RAHF.
@@ Appendix B Implementation Details
B.1 Training Setups
All baselines and our models were trained using Anthropic’s Helpful and Harmless datasetBai et al. (2022) fine-tuned model as the base model. During the supervised training of the base model, we calculated the loss for both prompts and responses. Specifically, we performed full parameter fine-tuning for three epochs with a learning rate of $52e-$ 5.
[…]
B.2 Evaluation Setups
For all methods, we employ greedy decoding during generation on the benchmarks. To avoid the issue of repetition, we set the repetition penalty to 1.2.
For the Open LLM Leaderboard, we utilized the Eleuther AI Language Model Evaluation Harness libraryGao et al. (2023) to assess language models trained using different methods. Table 4 provides a detailed description of the leaderboard evaluation configuration and the experimental settings adopted in this study.
  Datasets # few-shot Metric
Arc $25$ acc_norm
TruthfulQA 00 mc2
Winogrande $5$ acc
GSM8k $5$ acc
HellaSwag $10$ acc_norm
MMLU $5$ acc
Table 4: For each dataset used in the evaluation on the Open LLM Leaderboard, we detail the quantity of few-shot samples utilized and the specific metric employed for evaluation.
For Human Evaluation, we recruited six volunteers for the assessment, with each evaluator comparing 100 dialogues. Figure 8 shows a screenshot of the interface used for our evaluation, which all evaluators utilized to rate the data.
B.3 Experimental Details
In this section, we present the experimental details and hyperparameters of the baselines we compare with and our proposed methods.
Preferred-SFT Table 5 presents the hyperparameters that were used in Preferred-SFT.
  Hyperparameter Value
Learning Rate $52e-$ 5
Epochs $2$
Batch Size $128$
Micro Batch Size $2$
Optimizer Adamw
LR Scheduler Type Cosine
Rarmup Ratio $0.1$
Table 5: Hyperparameters used for Preferred-SFT.
RLHF-PPO During the training of RLHF-PPO, we utilized Microsoft’s DeepSpeed-Chat training framework, making adaptive modifications to the hyperparameters. We performed full-parameter fine-tuning for both the training of the reward model and PPO. Table 6 presents the hyperparameters for reward model training, while Table 7 presents the key parameters for PPO.
  Hyperparameter Value
Learning Rate $69.65e-$ 6
Epochs $3$
Optimizer Adam
Training Batch Size $32$
Weight Decay $0.1$
Warmup Steps 00
LR Scheduler Type cosine